# How to Tell When a Result Will Replicate: Significance and Replication in Distributional Null Hypothesis Tests

### v2

@@ 2 Significance and replication for the $t$ distribution
with unknown mean $\mu$ and unknown variance $\sigma^{2}$ (the standard assumption of normally distributed error), so that the sample mean $\overline{X}$ follows the distribution $\mathcal{N}(\mu,\sigma^{2}/N)$. Given sample variance $S^{2}$ and degrees of freedom $\nu$, the variance in $\overline{X}$ is estimated by $S^{2}/N$ (the sample variance of the mean) and the standard approach assesses significance against the point-form null hypothesis $\mu=0$ via a $t$-test.
@@ 2.1 Significance for a given variance ratio
is normally distributed with mean 00 and variance $1$. Letting
@@
as before, we thus see that
$\frac{\frac{\overline{X}}{(\sigma/\sqrt{N})\sqrt{1+bN}}}{\sqrt{(\frac{\nu\ S^{2}}{\sigma^{2}})/\nu}}=\frac{t}{\sqrt{1+bN}}$
follows a $t$-distribution with $\nu$ degrees of freedom, and so the statistical significance of a result $t$ in a two-sided test relative to the variance ratio $b$ is
$p_{sig}(t|b)=2T_{\nu}\left(-|t|/\sqrt{1+bN}\right)$ (1)
@@ 2.3 Significance and effect size
The size of an effect in a $t$-test is typically estimated in units of sample standard deviation, giving an effect-size estimate of $d=\overline{X}/S$ (an estimate of the true effect $\delta=\mu/\sigma$). Since in the point-form approach every experiment is expected to achieve significance for large enough $N$, effect size is necessarily independent of significance in this approach: the point-form significance of a result tells us nothing about its effect size. Effect size is thus considered separately from significance, with researchers proposing various criteria to distinguish between strong, weak, and negligible effects (Cohen,, 1992; Fritz et al.,, 2012; Kelley and Preacher,, 2012; Cohen,, 2013) alongside various lower bounds on acceptable effect sizes in terms of, for example, experimental ‘crud’ (Meehl, 1990b, ; Meehl, 1990a, ; Orben and Lakens,, 2020).
@@ 2.4 Power, $\beta$ and type $2$ error rate
To estimate $\beta$ for tests relative to a point-form null hypothesis, we let $T_{\nu}(x;\theta)$ represent the cumulative probability from $\infty$- ∞ to $x$ of the non-central $T$ distribution with degrees of freedom $\nu$ and non-centrality parameter $\theta$, and assume that $X\sim\mathcal{N}\left(\delta\sigma,\sigma^{2}\right)$ for some effect size $\delta\neq 0$ (so the null is false). In this situation the variable $t=(\overline{X}/S)\sqrt{N}$ follows a non-central $t$-distribution with non-centrality parameter $\delta\sqrt{N}$. The probability, in the point-form approach, of getting a non-significant result in a two-sided test with critical value $t_{crit}$, given that effect $\delta$, is then
@@
follows a non-central $t$-distribution with $\nu$ degrees of freedom and non-centrality parameter
@@ 2.5 Estimating the probability of replication
Because this distributional approach explicitly represents the variation in results across experiments (via the distributional null hypothesis), it allows us to express the probability of obtaining a statistically significant result at level $\alpha$ in a replication of our original experiment. In this section we give estimates for the replication probability, $p_{rep}$, for a given result based first on a fixed value $b$, and then based on a sample estimate $\hat{b}$. In subsequent sections we give simple ‘closed form’ approximations to $p_{sig}$ and $p_{rep}$ in terms of the $T$ distribution (for significance) and the Normal and inverse $\chi^{2}$ distributions (for replication). We also give generic bounds on $p_{sig}$ and $p_{rep}$ that can be used when no estimate for the variance ratio $\hat{b}$ is available.
For some value of $b$ we assume a replication experiment measuring the same $X$ as the original, with sample mean $\overline{X}_{r}\sim\mathcal{N}(\mu_{r},\,\sigma_{r}^{2}/N_{r})$, where $\mu_{r}$ and $\sigma^{2}_{r}$ are the unknown mean and unknown variance of $X$ in the replication, and $N_{r}$ and $\nu_{r}$ the number of measurements and the degrees of freedom. We take $S_{r}^{2}$ to represent the sample variance of $X$ in the replication, and relate within-experiment variance in the original and replication via the ratio $c=\sigma_{r}^{2}/\sigma^{2}$.
@@
be the test statistic, a result $t_{r}$ will count as a replication of our original result $t$ if it is significant at the same level $\alpha$, and in the same direction, as $t$. In other words, if our original result $t$ was significant in a left-tailed test ($t<-t_{crit}\sqrt{1+bN}$) then it is replicated if $t_{r}<-t_{crit_{r}}\sqrt{1+bN_{r}}$, while if our original result was significant in a right-tailed test ($t>t_{crit}\sqrt{1+bN}$) then it is replicated if $t_{r}>t_{crit_{r}}\sqrt{1+bN_{r}}$.
@@
has a $t$-distribution with $\nu_{r}$ degrees of freedom.
@@
will follow a $t$-distribution with $\nu_{r}$ degrees of freedom.
@@ 3 Estimating significance and replication
The previous section derived distributional null expressions for significance, power and replication assuming known variance ratio $b$ or bounds on that ratio. These expressions are both computationally complex, and depend on this unknown value. In this section we derive sample estimators for $b$, and give simple and easily used closed-form estimates for significance and replication in terms of the standard $T$ distribution. We also describe extensions of this approach to different forms of $t$-test and to tests of regression, linear correlation, and contingency.
@@ 3.1 Sample Estimators for $b$
Assuming that sample variances $S^{2}$ and $S_{r}^{2}$ in the two experiments are both estimates of the same underlying within-experiment variance $\sigma^{2}$, then since these two sample variances both follow a $\chi^{2}$ distribution with the same mean and with degrees of freedom $\nu$ and $\nu_{r}$, their ratio $c$ follows an $F$ distribution $F_{\nu,\nu_{r}}$. Similarly taking $\hat{b}=S_{0}^{2}/S^{2}$ as our estimate for the ratio $b$ we see that the variable $p_{rep}$ is distributed as a function of the joint distribution of $b$ and $c$. Since we don’t know the sample variance $S_{r}^{2}$ (our replication experiment has not yet been carried out) we simply set $\hat{c}=1$ and the expected value for $p_{rep}$ given estimate $\hat{b}$ is
@@ 3.3 Extension to related statistical tests
We’ve presented distributional significance testing here in terms of a generalised $t$-test, and so this approach applies to each distinct form of that test. To apply this model to a one-sample $t$-test on a variable $X$ with sample size $N$, we set $\nu=N-1$ and take $S^{2}$ to be the sample variance of that variable. To apply to a paired $t$-test on two variables (assuming $N$ pairs) we take the $X$ to represent the difference in pairs and $\overline{X}$ and $S^{2}$ to be the average difference between pairs and the variance of that difference, and $\nu=N-1$. To apply to an unpaired or independent $t$-test on two variables (with $N$ measurements in total), we take $\overline{X}$ to be the difference between the means of each variable, $N$ to be the sample size, and $S^{2}$ to be the pooled sample variance with degrees of freedom $\nu=N-2$. (Note that we implicitly assume an equal number of samples for the two variables in an unpaired test. This is because the probability of replication of a given result in a future experiment on two unpaired variables will depend on the sample sizes for those two variables in that experiment, which we cannot know. Assuming that these sample sizes vary randomly the expected sample size for each variable will be $N/2$; and so we use the calculated replication probability assuming equal sample sizes as our estimate of the probability of replication.)
The approach also applies to tests of linear regression, correlation and contingency, which can all be related to the $t$-test. Applying the distributional approach to tests of independence for correlation and linear regression gives $p_{sig}$ and $p_{rep}$ expressions as above, albeit calculated with sample slope $M$ replacing sample mean $\overline{X}$, predictor sum-of-squares (which we write as $Q$) replacing sample size $N$, and mean squared error of the response estimate replacing sample variance. This is because, with these replacements, such tests of independence are simply $t$-tests (see Appendix $2$ for a derivation of these results). Note that in these correlation tests the sum of squares of the predictor variable is the continuous-valued analog of the discrete sample size $N$ in a $t$-test, and is treated as a controlled variable: where sample size gives the number of discrete measurements made in a $t$-test, predictor sum-of-squares gives the range of continuous measurements in a test of correlation. This means that the significance and replication probability of these tests depends on the choice of predictor and response variables: an estimate for the probability of replication of a given result in an experiment where $X$ is the predictor (and so assumed to be controlled in replications) will necessarily be different from the probability of replication of that result when $Y$ is the predictor and the $Y$ sum of squares is assumed to be controlled in replications (because the two replication experiments themselves are different).
@@ 3.4 Comparing with point-form approaches to replication
Our distributional null hypothesis expression for $p_{rep}$ resolves the problems in Killeen’s approach. Our $p_{rep}$ is not simply a function of the point-form $p$ and is an estimate for replication in a particular experiment (in that it depends on the ratio of between-experiment to within experiment variance $b$, the ratio of within-experiment variance in the original experiment and the replication $c$, and the sample sizes of the original and replication experiments). Our $p_{rep}$ is specifically based on a distribution for the null hypothesis, does not assume all effect sizes are equally likely (the larger an effect size the less likely it is under this distribution), and does not assume the null hypothesis is false. Finally, since our result is based on well-known properties of the Normal and $t$ distributions, errors in the derivation are (hopefully) unlikely, and the associated claims have a clear interpretation.
@@ 4.1 Methods
Task
Total test $S_{0}^{2}$ $\hat{b}$ same direction as original
@@
1
Allowed/forbidden
4934 $\chi^{2}$ 0.005 0.041 1.00 1.00 $10^{-325}$ $10^{-10}$
2
Anchoring - babies born
4481 t $6.4\times 10^{6}$ 0.06 1.00 1.00 $10^{-325}$ $10^{-4}$
3
Anchoring - Mt. Everest
4575 t $9.3\times 10^{6}$ 0.11 1.00 1.00 $10^{-325}$ 0.0004
4
Anchoring - Chicago
4302 t $8.1\times 10^{10}$ 0.077 1.00 1.00 $10^{-325}$ 0.001
5
Anchoring - distance to NYC
4422 t $5.0\times 10^{4}$ 0.05 1.00 0.76 $10^{-266}$ 0.01
6
Relation between Implicit and Explicit math attitudes
4383 r 0.080 0.12 0.88 0.52 $10^{-148}$ 0.01
7
Retrospective gambler fallacy
4690 t 0.640 0.12 0.88 0.00 $10^{-98}$ 0.15
8
Low vs High category scales
4664 $\chi^{2}$ 0.004 0.077 0.80 0.28 $10^{-57}$ 0.01
9
Gain vs. loss framing
4925 $\chi^{2}$ 0.007 0.029 0.84 0.68 $10^{-91}$ 0.001
10
Sex differences in implicit math attitudes
4558 t 0.010 0.05 0.76 0.00 $10^{-56}$ 0.15
11
quote attribution
4964 t 0.260 0.063 0.56 0.00 $10^{-36}$ 0.23
12
sunk costs
4967 t 0.110 0.032 0.36 0.00 $10^{-19}$ 0.19
13
norm of reciprocity
4928 $\chi^{2}$ 0.008 0.04 0.44 0.20 $10^{-28}$ 0.07
14
imagined contact
4973 t 0.110 0.034 0.16 0.00 $10^{-7}$ 0.34
15
Flag priming
4896 t 0.012 0.017 0.00 0.00 0.20 0.45
16
Currency priming
4970 t 0.016 0.021 0.00 0.00 0.40 0.49
@@ 4.3 Comparing point-form and distributional significance
As an initial check we used numerical integration methods to calculate, for each experiment, the theoretical value $p_{sig}(t|\hat{b})$ given that experiment’s $t$ value and the $\hat{b}$ estimate for that task, and compared against the closed-form expression $p_{sig}$ for that experiment. $p_{sig}p(t|\hat{b})$ and $p_{sig}$ values were almost identical (correlation $r=0.99$), supporting the closed-form approximations, and so we used only the closed-form approximation $p_{sig}$ in our analysis.
@@
Figure 2: (Left) Scatterplot of $\log_{10}$ distributional $p_{sig}$ against $\log_{10}$ point-form $p$ for all experiments in all tasks in the Many Labs 1 replication dataset for US sites (note the difference in scales). Lines indicate statistical significance at levels $\alpha=0.05$ (dotted), $\alpha=0.025$ (dashed) and $\alpha=0.01$ (solid). (Right) Analogous scatterplot of $\log_{10}$ distributional $p_{sig}$ against $\log_{10}$ Bayes Factor $BF_{10}$. Horizontal lines indicate Bayes Factor levels of $3$ (dotted: $BF_{10}\geq 3$ being typically interpreted as moderate evidence against the null), $10$ (dashed: strong evidence against the null) and $30$ (solid:extremely strong evidence against the null). Points for $9$ experiments, all with $p<10^{-40}$ ($BF_{10}>10^{40}$) and ranging as low as $p<10^{-325}$ ($BF_{10}>10^{325}$), are not shown.

### v3

@@ 2 Significance and replication for the $t$ distribution
with unknown mean $\mu$ and unknown variance $\sigma^{2}$ (the standard assumption of normally distributed error), so that the sample mean $\overline{X}$ follows the distribution $\mathcal{N}(\mu,\sigma^{2}/N)$. Given sample variance $S^{2}$ and degrees of freedom $\nu$, the variance in $\overline{X}$ is estimated by $S^{2}/N$ (the sample variance of the mean) and the standard approach assesses significance against the point-form null hypothesis $\mu=0$ via a $t$ test.
@@ 2.1 Significance for a given variance ratio
is normally distributed with mean 00 and variance $1$. Letting $S^{2}$ be our sample variance (our estimate of $\sigma^{2}$) by definition
$\frac{S^{2}}{\sigma^{2}}$
follows a $\chi_{\nu}^{2}$ distribution (where $\nu$ is the degrees of freedom $S^{2}$). Letting
@@
as before and recalling that the ratio of a standard Normal variable to the square root of a $\chi_{\nu}^{2}$ variable follows a $t$ distribution with $\nu$ degrees of freedom, we see that the statistic
$\frac{\frac{\overline{X}}{(\sigma/\sqrt{N})\sqrt{1+bN}}}{\sqrt{\frac{S^{2}}{\sigma^{2}}}}=\frac{t}{\sqrt{1+bN}}$
is $t$ distributed. This means that the statistical significance of a given result $t$ in a two-sided test relative to the variance ratio $b$ is
$p_{sig}(t|b)=2T_{\nu}\left(\frac{-|t|}{\sqrt{1+bN}}\right)$ (1)
@@ 2.3 Significance and effect size
The size of an effect in a $t$ test is typically estimated in units of sample standard deviation, giving an effect-size estimate of $d=\overline{X}/S$ (an estimate of the true effect $\delta=\mu/\sigma$). Since in the point-form approach every experiment is expected to achieve significance for large enough $N$, effect size is necessarily independent of significance in this approach: the point-form significance of a result tells us nothing about its effect size. Effect size is thus considered separately from significance, with researchers proposing various criteria to distinguish between strong, weak, and negligible effects (Cohen,, 1992; Fritz et al.,, 2012; Kelley and Preacher,, 2012; Cohen,, 2013) alongside various lower bounds on acceptable effect sizes in terms of, for example, experimental ‘crud’ (Meehl, 1990b, ; Meehl, 1990a, ; Orben and Lakens,, 2020).
@@ 2.4 Power, $\beta$ and type $2$ error rate
To estimate $\beta$ for tests relative to a point-form null hypothesis, we let $T_{\nu}(x;\theta)$ represent the cumulative probability from $\infty$- ∞ to $x$ of the non-central $t$ distribution with degrees of freedom $\nu$ and non-centrality parameter $\theta$, and assume that $X\sim\mathcal{N}\left(\delta\sigma,\sigma^{2}\right)$ for some effect size $\delta\neq 0$ (so the null is false). In this situation the variable $t=(\overline{X}/S)\sqrt{N}$ follows a non-central $t$ distribution with non-centrality parameter $\delta\sqrt{N}$. The probability, in the point-form approach, of getting a non-significant result in a two-sided test with critical value $t_{crit}$, given that effect $\delta$, is then
@@
follows a non-central $t$ distribution with $\nu$ degrees of freedom and non-centrality parameter
@@ 2.5 Estimating the probability of replication
Because this distributional approach explicitly represents the variation in results across experiments (via the distributional null hypothesis), it allows us to express the probability of obtaining a statistically significant result at level $\alpha$ in a replication of our original experiment. In this section we give estimates for the replication probability, $p_{rep}$, for a given result based first on a fixed value $b$, and then based on a sample estimate $\hat{b}$.
For some value of $b$ we assume a replication experiment designed to measure the same $X$ as the original (but with some between-experiment variation), with sample mean for the replication experiment following the distribution $\overline{X}_{r}\sim\mathcal{N}(\mu_{r},\,\sigma_{r}^{2}/N_{r})$, where $\mu_{r}$ and $\sigma^{2}_{r}$ are the unknown mean and unknown variance of $X$ in the replication, and $N_{r}$ and $\nu_{r}$ the number of measurements and the degrees of freedom. We take $S_{r}^{2}$ to represent the sample variance of $X$ in the replication, and relate within-experiment variance in the original and replication via the ratio $c=\sigma_{r}^{2}/\sigma^{2}$.
@@
the experimental result, a result $t_{r}$ will count as a replication of our original result $t$ if both $t$ and $t_{r}$ are significant at the same level $\alpha$, and in the same direction, relative to our distributional null hypothesis. In other words, if our original result $t$ was significant in a left-tailed test ($t<-t_{crit}\sqrt{1+bN}$) then it is replicated if $t_{r}<-t_{crit_{r}}\sqrt{1+bN_{r}}$, while if our original result was significant in a right-tailed test ($t>t_{crit}\sqrt{1+bN}$) then it is replicated if $t_{r}>t_{crit_{r}}\sqrt{1+bN_{r}}$.
@@
has a $t$ distribution with $\nu_{r}$ degrees of freedom.
@@
will follow a $t$ distribution with $\nu_{r}$ degrees of freedom.
@@ 3 Estimating significance and replication
The previous section derived distributional null expressions for significance, power and replication assuming known variance ratio $b$ or bounds on that ratio. These expressions are both computationally complex, and depend on this unknown value. In this section we derive sample estimators for $b$, and give simple and easily used closed-form estimates for significance and replication in terms of the standard $t$ distribution. We also describe extensions of this approach to different forms of $t$ test and to tests of regression, linear correlation, and contingency.
@@ 3.1 Sample Estimators for $b$
To consider the probability of replication in a new experiment, we assume that sample variances $S^{2}$ and $S_{r}^{2}$ in the two experiments are both estimates of the same underlying within-experiment variance $\sigma^{2}$ as before. Then since these two sample variances both follow a $\chi^{2}$ distribution with the same mean and with degrees of freedom $\nu$ and $\nu_{r}$, their ratio $c$ follows an $F$ distribution $F_{\nu,\nu_{r}}$. Similarly taking $\hat{b}=S_{0}^{2}/S^{2}$ as our estimate for the ratio $b$ we see that the variable $p_{rep}$ is distributed as a function of the joint distribution of $b$ and $c$. Since we don’t know the sample variance $S_{r}^{2}$ (our replication experiment has not yet been carried out) we simply set $\hat{c}=1$ and the expected value for $p_{rep}$ given estimate $\hat{b}$ is
@@ 3.3 Extension to related statistical tests
We’ve presented distributional significance testing here in terms of a generalised $t$ test, and so this approach applies to each distinct form of that test. To apply this model to a one-sample $t$ test on a variable $X$ with sample size $N$, we set $\nu=N-1$ and take $S^{2}$ to be the sample variance of that variable. To apply to a paired $t$ test on two variables (assuming $N$ pairs) we take the $X$ to represent the difference in pairs and $\overline{X}$ and $S^{2}$ to be the average difference between pairs and the variance of that difference, and $\nu=N-1$. To apply to an unpaired or independent $t$ test on two variables (with $N$ measurements in total), we take $\overline{X}$ to be the difference between the means of each variable, $N$ to be the sample size, and $S^{2}$ to be the pooled sample variance with degrees of freedom $\nu=N-2$. (Note that we implicitly assume an equal number of samples for the two variables in an unpaired test. This is because the probability of replication of a given result in a future experiment on two unpaired variables will depend on the sample sizes for those two variables in that experiment, which we cannot know. Assuming that these sample sizes vary randomly the expected sample size for each variable will be $N/2$; and so we use the calculated replication probability assuming equal sample sizes as our estimate of the probability of replication.)
The approach also applies to tests of linear regression, correlation and contingency, which can all be related to the $t$ test. Applying the distributional approach to tests of independence for correlation and linear regression gives $p_{sig}$ and $p_{rep}$ expressions as above, albeit calculated with sample slope $M$ replacing sample mean $\overline{X}$, predictor sum-of-squares (which we write as $Q$) replacing sample size $N$, and mean squared error of the response estimate replacing sample variance. This is because, with these replacements, such tests of independence are simply $t$ tests (see Appendix $2$ for a derivation of these results). Note that in these correlation tests the sum of squares of the predictor variable is the continuous-valued analog of the discrete sample size $N$ in a $t$ test, and is treated as a controlled variable: where sample size gives the number of discrete measurements made in a $t$ test, predictor sum-of-squares gives the range of continuous measurements in a test of correlation. This means that the significance and replication probability of these tests depends on the choice of predictor and response variables: an estimate for the probability of replication of a given result in an experiment where $X$ is the predictor (and so assumed to be controlled in replications) will necessarily be different from the probability of replication of that result when $Y$ is the predictor and the $Y$ sum of squares is assumed to be controlled in replications (because the two replication experiments themselves are different).
@@ 3.4 Comparing with point-form approaches to replication
Our distributional null hypothesis expression for $p_{rep}$ resolves the problems in Killeen’s approach. Our $p_{rep}$ is not simply a function of the point-form $p$ and is an estimate for replication in a particular experiment (in that it depends on the ratio of between-experiment to within experiment variance $b$, the ratio of within-experiment variance in the original experiment and the replication $c$, and the sample sizes of the original and replication experiments). Our $p_{rep}$ is specifically based on a distribution for the null hypothesis, does not assume all effect sizes are equally likely (the larger an effect size the less likely it is under this distribution), and does not assume the null hypothesis is false. Finally, since our result is based on well-known properties of the Normal and $t$ distributions, errors in the derivation are (hopefully) avoided, and the associated claims have a clear interpretation.
@@ 4.1 Methods
Task Total test $S_{0}^{2}$ $\hat{b}$ same direction as original
@@
1 Allowed/forbidden 4934 $\chi^{2}$ 0.005 0.041 1.00 1.00 $10^{-325}$ $10^{-10}$
2 Anchoring - babies born 4481 $t$ $6.4\times 10^{6}$ 0.06 1.00 1.00 $10^{-325}$ $10^{-4}$
3 Anchoring - Mt. Everest 4575 $t$ $9.3\times 10^{6}$ 0.11 1.00 1.00 $10^{-325}$ 0.0004
4 Anchoring - Chicago 4302 $t$ $8.1\times 10^{10}$ 0.077 1.00 1.00 $10^{-325}$ 0.001
5 Anchoring - distance to NYC 4422 $t$ $5.0\times 10^{4}$ 0.05 1.00 0.76 $10^{-266}$ 0.01
6 Relation between Implicit and Explicit math attitudes 4383 $r$ 0.080 0.12 0.88 0.52 $10^{-148}$ 0.01
7 Retrospective gambler fallacy 4690 $t$ 0.640 0.12 0.88 0.00 $10^{-98}$ 0.15
8 Low vs High category scales 4664 $\chi^{2}$ 0.004 0.077 0.80 0.28 $10^{-57}$ 0.01
9 Gain vs. loss framing 4925 $\chi^{2}$ 0.007 0.029 0.84 0.68 $10^{-91}$ 0.001
10 Sex differences in implicit math attitudes 4558 $t$ 0.010 0.05 0.76 0.00 $10^{-56}$ 0.15
11 quote attribution 4964 $t$ 0.260 0.063 0.56 0.00 $10^{-36}$ 0.23
12 sunk costs 4967 $t$ 0.110 0.032 0.36 0.00 $10^{-19}$ 0.19
13 norm of reciprocity 4928 $\chi^{2}$ 0.008 0.04 0.44 0.20 $10^{-28}$ 0.07
14 imagined contact 4973 $t$ 0.110 0.034 0.16 0.00 $10^{-7}$ 0.34
15 Flag priming 4896 $t$ 0.012 0.017 0.00 0.00 0.20 0.45
16 Currency priming 4970 $t$ 0.016 0.021 0.00 0.00 0.40 0.49
@@ 4.3 Comparing point-form and distributional significance
As an initial check we used numerical integration methods to calculate, for each experiment, the theoretical value $p_{sig}(t|\hat{b})$ given that experiment’s $t$ value and the $\hat{b}$ estimate for that task, and compared against the closed-form expression $p_{sig}$ for that experiment. $p_{sig}(t|\hat{b})$ and $p_{sig}$ values were almost identical (correlation $r=0.99$), supporting the closed-form approximations, and so we used only the closed-form approximation $p_{sig}$ in our analysis.
@@
Figure 2: (Left) Scatterplot of $\log_{10}$ distributional $p_{sig}$ against $\log_{10}$ point-form $p$ for all experiments in all tasks in the Many Labs 1 replication dataset for US sites (note the difference in scales). Lines indicate statistical significance at levels $\alpha=0.05$ (dotted), $\alpha=0.025$ (dashed) and $\alpha=0.01$ (solid). (Right) Analogous scatterplot of $\log_{10}$ distributional $p_{sig}$ against $\log_{10}$ Bayes Factor $BF_{10}$. Horizontal lines indicate Bayes Factor levels of $3$ (dotted: $BF_{10}\geq 3$ being typically interpreted as moderate evidence against the null), $10$ (dashed: strong evidence against the null) and $30$ (solid:extremely strong evidence against the null). Points for $9$ experiments, all with $p<10^{-40}$ ($BF_{10}>10^{40}$) and ranging as low as $p<10^{-325}$ ($BF_{10}>10^{325}$), are not shown. Note that $\log_{10}p$ and $\log_{10}BF$ values are identical modulo difference in scale (a $r=0.9999$ correlation between $\log_{10}p$ and $\log_{10}BF$ across experiments).